OpenStudy (anonymous):

7.2-11). Multiply the following expression. (x^2-x-2) -7x ______ x+1

OpenStudy (anonymous):

\[-\frac{7 x ((x-2) (x+1))}{x+1}=-7 (x-2) x=14 x-7 x^2 \]